Status Report

Hubble Space Telescope Daily Report #2942 23 Aug 2001

By SpaceRef Editor
August 23, 2001
Filed under , ,

The Space Telescope Imaging Spectrograph (CCD) was used to measure
D/H in gas clouds seen against background QSOs. This provides a unique and
fundamental cosmological probe of the baryonic density parameter. To date,
most of the QSO absorption line D/H constraints have been derived at high
redshift. In this proposal the proposers investigate a lower redshift
range, because there are two potentially crucial advantages to be gained by
carrying out D/H measurements at intermediate redshift. First, the
background of HI forest lines is lower and contamination of the DI feature
is significantly reduced. Second, a far larger sample of very bright QSOs
exists. There were no reported problems.


2.1 Guide Star Acquisitions:

Scheduled Acquisitions: 6

Successful: 6

As described in HSTAR 8314, the acquisition at 234/155929Z
defaulted to fine lock backup on FGS-3 only when the scan step limit was
exceeded on FGS-2. The proposals detailed in 1.9 and 1.10 may have been
Scheduled Re-acquisitions: 9

Successful: 9

2.2 FHST Updates:

Scheduled: 12

Successful: 12

2.3 Operations Notes:

The SSR EDAC error counter was cleared twice per ROP SR-1A.

Flight software 2.0B EEPROM and 2.0C RAM installation occurred
successfully at 234/1330Z. This installation, together with related
installations performed yesterday and another installation tomorrow,
changes the post-SM3B Diode Bus Controller commanding and fixes the
Magnetic Torquer echo check timing.


Installation of 2.0C EEPROM flight software today.

Continuation of normal science observations and calibrations.

SpaceRef staff editor.